Suppose you have the following information about a fictitious economy. Assume there are no taxes in this economy.

Disposable Income and Consumption

Disposable Income (dollars) Consumption (dollars)
$0 $9,000
8,500 15,500
18,500 22,000
28,500 28,500
38,500 35,000
48,500 41,500

Instructions: In parts a and c, enter your answers as a whole number. In part b, round your answers to two decimal places.

a. What is the equilibrium level of consumption?

     $_________

b. What is the MPC and MPS for this economy?

     MPC: _________

     MPS: _________

c. What is the value of autonomous consumption for this economy?

     $ _________
Answer #1

A) $ 28,500

B) MPC = ( 15,500 - 9,000) / 8,500 = 0.76

MPS = 1 - 0.76 = 0.24

C) $ 9,000